OpenStudy (anonymous):

Find the equation of the line whose x-intercept is 8 and whose y-intercept is -3, in standard form.

OpenStudy (anonymous):

it's mean you given points: (8,0) & (0,-3) your slope is m=3/8

OpenStudy (anonymous):

your equation is : y=3/8 x -3

OpenStudy (anonymous):

or: 8y=3x-24 8y-3x=-24 8y-3x+24=0